Question as given - the first four terms in the series sn=13-3n
I've just gone back to college after a long long sabbatical. This professor doesn't actually work/do problems on the board. We're to read the book and figure it out. I've made it through chapter 11.2 and I just cannot figure this out on my own.

Answer by macston(5194)

.
Solve the equation for n=1 to 4 (first 4):
s(1)=13-3(1)=13-3=10
s(2)=13-3(2)=13-6=7
s(3)=13-3(3)=13-9=4
s(4)=13-3(4)=13-12=1